# Sustainable and Precision Agriculture with the Internet of Everything (IoE)

### v2

@@ II-A Overview
  • Data reliability: Ensuring data reliability in IoT systems is vital. Issues such as “fail-dirty” phenomenon [36], where a sensor malfunctions and sends erroneous data can have detrimental consequences. This particular issue is difficult to diagnose as the sensor appears to be in normal operating condition. Addressing this issue requires strategies like redundancy and the use of machine learning (ML) algorithms.
@@
II-B Categorization of Application Areas
@@
III Advancing smart Agricultural through Nanotechnology and Nano Communication
IoT is mainly about connecting tangible devices to the Internet. In contrast, IoE broadens this concept by integrating computing devices, inanimate and living entities, people, data, and operational processes into the network. Meanwhile, Internet of Nano things (IoNT) takes this idea even further, venturing into the realm of nano-scale technology for more advanced connectivity [57]. This section explores the contributions of nanotechnology, molecular communication (MC), Internet of Nano Things (IoNT) and Internet of Bio-Nano Things (IoBNT) in the realization of smart agriculture, with a summary of applications, challenges and future research directions presented in Table I.
III-A Impact of Nanomaterials & Nanoparticles in Agriculture
@@
III-B Enhancing Smart Agriculture through Molecular Communication
@@
III-C Internet of Nano Things (IoNT) for Agriculture
@@
III-D Role of Internet of Bio-Nano Things (IoBNT) in Agricultural Innovation
IoNT has revolutionized the technological landscape through the advancements in nanomaterials like graphene, facilitating the interconnection of nanoscale devices. Despite its marvellous benefits, IoNT could pose harmful effects on health or contribute to pollution. This challenge has led to the integration of biological entities and nanotechnology giving rise to the Internet of Bio-Nano Things (IoBNT) [70], [79]. Along with other critical applications such as healthcare, IoBNT is poised to play a vital role in agriculture with biosensors and actuators to function inside or near plant body.
@@ III-D Role of Internet of Bio-Nano Things (IoBNT) in Agricultural Innovation
TABLE I: Applications, Challenges and Research Directions of Nanotechnology and Nanomaterials in Smart Agriculture
@@
IV Harnessing Engineered and Natural Phages to Augment Animal and Plant Health
Bacteriophages, also known as phages, are viruses that specifically infect bacteria. They are ubiquitous and offer a wide range of applications. With the emergence of designer phages in microbiome engineering, we are at a crucial juncture poised to unravel a vast array of possibilities. These meticulously crafted biological agents hold great promise in revolutionizing the health of plants and animals, markedly improving agricultural productivity, and offering novel approaches to address the daunting issue of antibiotic resistant bacteria. In this section, the role of phages in plant and animal health, techniques & technologies used, and some future perspectives will be discussed. Additionally, an overview of their applications, the challenges encountered, and directions for future research are summarized in Table II.
IV-A Introduction to Phage Therapy
@@
IV-B Phage Therapy for the Health of Plants & Animals
@@
IV-C Techniques and Technologies used in Phages
@@
TABLE II: Applications, Challenges and Research Directions of Phages in Smart Agriculture
@@
IV-D Future Directions & Research Opportunities
@@
V Internet of Fungus for Modern Agriculture
Humans have developed various platforms and methods to facilitate information exchange and provide mutual assistance. Similarly, in the natural world all living organism, from animals to plants, engage in communication. Among plants this communication occurs due to the mycelial (fungus) network, which interconnects plants and helps information and nutrients exchange between them. The internet of fungus is vital in modern agriculture, particularly for its role in the improvement of plant health, plant security, nutrients delivery and overall crops productivity. In this section, we will explore the concept of “Internet of Fungus”, examine their communication mechanisms, applications within agriculture and some potential future research directions and opportunities. In addition, a succinct summary is provided in Table III.
V-A Introduction to the Internet of Fungus
@@
V-B Communication Mechanism within Fungal Networks
@@ V-C Applications in Agriculture
TABLE III: Applications, Challenges and Research Directions of Internet of Fungus in Smart Agriculture
@@
V-D Future Directions & Research Opportunities
The “internet of fungus” is an innovative paradigm that holds immense potential in revolutionizing the agricultural sector. It can improve crop & soil health, boost productivity and provide a sustainable ecosystem. This concept underscores the importance of investigating fungal networks to advance agricultural practices. Some prospective areas of research are mentioned below.
@@ V-D Future Directions & Research Opportunities
VI Internet of Energy Harvesting Things: Sustainable Power for Agriculture
In the IoT, ensuring sustainable power for a vast network of interconnected devices is a critical challenge. Conventional power sources, such as direct electrical connections or bulky batteries, are not viable for maintaining the multitude of small, ubiquitous smart objects continuously operational. Energy harvesting emerges as a vital solution, enabling these devices to operate autonomously and overcome power constraints. In this section, we will delve into the realm of energy harvesting for smart agriculture, including its techniques, current trends, and future directions & research opportunities. Furthermore, a summary of applications, challenges and future research directions is presented in Table IV.
VI-A Overview of Internet of Energy
VI-B Energy Harvesting Techniques
  1. 1.
    TABLE IV: Applications, Challenges and Research Directions of Internet of Energy in Smart Agriculture
@@ VI-B Energy Harvesting Techniques
VI-C Emerging Trends in Energy Harvesting
@@
VI-D Future Directions & Research Opportunities
@@
VII The Synergy of Internet of Vehicles, Internet of Drones & Internet of Space Things for Efficient Farming
The convergence of IoT, Internet of Vehicles (IoV), Internet of Drones (IoD) and Internet of Space Things (IoST) is anticipated to significantly transform the agricultural sector. This integration will provide farmers and researchers with real-time insights, enhanced resource management, and the facilitation of sustainable operations & practices. Specifically, internet connected tractors, harvesters and other agriculture machinery will enhance the productivity levels. Similarly, IoD and IoST frameworks are set to empower farmers to make data-driven and informed decisions. Fig. 7 depicts the synergistic impact of IoST, IoD, IoV, and IoEn on agriculture. In this section we will discuss the pivotal role of IoV, IoD and IoST for augmenting the productivity of agriculture sector. Additionally, some prospective future directions will also be explored, with a summary provided in Table V.
VII-A Internet of Vehicles: Optimizing Agricultural Operations
@@
VII-B Internet of Drones: A new Horizon in Pest Management and Crop Surveillance
@@
VII-C Internet of Space Things: Taking Agricultural Data to New Heights
TABLE V: Applications, Challenges and Research Directions of IoV, IoD & IoST in Smart Agriculture
@@ VII-C Internet of Space Things: Taking Agricultural Data to New Heights
VII-D Future Directions & Research Opportunities
@@
VIII The Nexus of 6G and Precision Agriculture
The Internet of Things (IoT), a vast network connecting a multitude of objects and devices, has been instrumental in fostering novel services and technological innovations. As the number of IoT devices continues to surge, there is an increasing demand for robust network infrastructure. Alongside the widespread adoption of IoT, several challenges emerge that require resolution. This interconnected ecosystem depends heavily on sophisticated communication technologies, such as 5G and the upcoming 6G network [186], which are crucial for enabling advanced hybrid and multi-cloud solutions. In this section we will discuss the critical role of 6G technology in enhanced connectivity for modern agriculture, key technological enablers of 6G and some prospective future research directions. In addition, a summary of applications, challenges and future research directions is presented in Table VI.
VIII-A Enhanced Connectivity for Modern Agriculture
@@
VIII-B 6G Technological Enablers for Agricultural Advancements
@@
    Figure 9: Key technologies for enabling 6G communication
@@
    TABLE VI: Applications, Challenges and Research Directions of 6G in Smart Agriculture
    6G
      High energy consumption due to high frequency bands [233].
      Expensive 6G infrastructure.
      Lack of legal & policy frameworks.
      Integrating AI with 6G for NW management in IoAT.
      Zero-Energy sensors for IoAT.
      THz spectroscopy for metal contaminant detection in food.
@@
VIII-C Future Directions & Research Opportunities
@@
IX Blockchain Technology for Smart & Sustainable Agriculture Sector
Blockchain technology is a digital system that enables electronic documentation, validation and verification thereby eliminating the need for intermediaries. It ensures data and information visibility for all participants, with all records being immutable that are impervious to tampering or erasure. The latest trends and improvements in blockchain technology offer some exciting features that are well suitable for agriculture. This section provides an overview of the blockchain technology, explores its applications & challenges within the agricultural sector, and highlights potential future research directions and opportunities, with a summary presented in Table VII.
IX-A Overview of Blockchain Technology
@@
Blockchain’s contributions to high security, enhanced trust and improved efficiency enable its application beyond mere financial transactions. It extends into areas such as supply chain management, voting systems, and data protection, thus reshaping our digital interactions with a new level of transparency and immutability [253].
IX-B Blockchain Technology in Agriculture
@@
    Centralized storage, typical of IoT-based systems, makes data susceptible to loss and manipulation. A distributed database can mitigate this issue by facilitating a decentralized storage.
@@
    Blockchain, in this context, provides a secure, scalable and robust framework for IoT and ICTs in the realm of smart farming. Several smart farming models based on blockchain & IoT have been successfully implemented. In [265] authors introduced an Ethereum blockchain network designed for automating water control in irrigation systems. A generalized blockchain based security architecture for smart sensing in IoT driven agriculture is presented in [266]. [267] explores smart contracts for enabling scalable data sharing in the context of IoT for smart agriculture. Comprehensive literature reviews on blockchain & IoT are provided in [268], [269], [270], [271], [272], underscoring the importance of these technologies in advancing precision agriculture.
@@
IX-C Future Directions & Research Opportunities
@@
TABLE VII: Applications, Challenges and Research Directions of Blockchain in Smart Agriculture
Blockchain Technology
Transparency & Security   Supply chain management.
  Secure data sharing.
  Agriculture finance.
  Interoperability across blockchain networks.
  High energy consumption by blockchains.
  Lack of regulatory policies.
  Addressing scalability issues in large scale deployments.
  Integrating energy harvesting to reduce costs.
  Tokenizing agricultural assets for new opportunities.
  Develop quantum blockchain hybrid models.
  Agricultural insurance to facilitate compensations for losses.
X Machine learning (ML) in Precision Agriculture
The incorporation of IoT into agriculture has significantly increased this sector’s efficiency and productivity through the goal of full process automation. There is a growing emphasis among researchers on the use of intelligent computational methods to further augment agricultural practices. Numerous ML algorithms have been developed and applied for purposes like monitoring, controlling, predicting, and smart decision making. ML applications span across various sectors including agriculture, education, healthcare, military, finance, transportation, marketing, manufacturing, and cybersecurity. In this section, we will explore the basics of ML, investigate its utilization in precision agriculture, and identify prospective directions for future research. In addition, a summary of applications, challenges, and future research directions is also presented in Table VIII.
@@ X Machine learning (ML) in Precision Agriculture
X-A Overview of Machine Learning
@@
X-B Machine Learning Applications in Precision Agriculture
@@
    TABLE VIII: Applications, Challenges and Research Directions of Machine Learning in Smart Agriculture
@@
X-C Future Directions & Research Opportunities
@@
XI Conclusion

### v3

@@ III-A Overview
  • Data reliability: Ensuring data reliability in IoT systems is vital. Issues such as “fail-dirty” phenomenon [36], where a sensor malfunctions and sends erroneous data can have detrimental consequences. This particular issue is difficult to diagnose as the sensor appears to be in normal operating condition. Addressing this issue requires strategies like redundancy and the use of ML algorithms.
@@
III-B Categorization of Application Areas
@@
IV Advancing smart Agricultural through Nanotechnology and Nano Communication
IoT is mainly about connecting tangible devices to the Internet. In contrast, IoE broadens this concept by integrating computing devices, inanimate and living entities, people, data, and operational processes into the network. Meanwhile, Internet of Nano things (IoNT) takes this idea even further, venturing into the realm of nano-scale technology for more advanced connectivity [56]. This section explores the contributions of nanotechnology, molecular communication (MC), IoNT and Internet of Bio-Nano Things (IoBNT) in the realization of smart agriculture, with a summary of applications, challenges and future research directions presented in Table V.
IV-A Impact of Nanomaterials & Nanoparticles in Agriculture
@@
IV-B Enhancing Smart Agriculture through Molecular Communication
@@
IV-C Internet of Nano Things for Agriculture
@@
IV-D Role of Internet of Bio-Nano Things in Agricultural Innovation
IoNT has revolutionized the technological landscape through the advancements in nanomaterials like graphene, facilitating the interconnection of nanoscale devices. Despite its marvellous benefits, IoNT could pose harmful effects on health or contribute to pollution. This challenge has led to the integration of biological entities and nanotechnology giving rise to the IoBNT [69], [78]. Along with other critical applications such as healthcare, IoBNT is poised to play a vital role in agriculture with biosensors and actuators to function inside or near plant body.
@@ IV-D Role of Internet of Bio-Nano Things in Agricultural Innovation
TABLE V: Applications, Challenges and Research Directions of Nanotechnology and Nanomaterials in Smart Agriculture
@@
V Harnessing Engineered and Natural Phages to Augment Animal and Plant Health
Bacteriophages, also known as phages, are viruses that specifically infect bacteria. They are ubiquitous and offer a wide range of applications. With the emergence of designer phages in microbiome engineering, we are at a crucial juncture poised to unravel a vast array of possibilities. These meticulously crafted biological agents hold great promise in revolutionizing the health of plants and animals, markedly improving agricultural productivity, and offering novel approaches to address the daunting issue of antibiotic resistant bacteria. In this section, the role of phages in plant and animal health, techniques & technologies used, and some future perspectives will be discussed. Additionally, an overview of their applications, the challenges encountered, and directions for future research are summarized in Table VI.
V-A Introduction to Phage Therapy
@@
V-B Phage Therapy for the Health of Plants & Animals
@@
V-C Techniques and Technologies used in Phages
@@
TABLE VI: Applications, Challenges and Research Directions of Phages in Smart Agriculture
@@
V-D Future Directions & Research Opportunities
@@
VI Internet of Fungus for Modern Agriculture
Humans have developed various platforms and methods to facilitate information exchange and provide mutual assistance. Similarly, in the natural world all living organism, from animals to plants, engage in communication. Among plants this communication occurs due to the mycelial (fungus) network, which interconnects plants and helps information and nutrients exchange between them. The internet of fungus (IoF) is vital in modern agriculture, particularly for its role in the improvement of plant health, plant security, nutrients delivery and overall crops productivity. In this section, we will explore the concept of IoF, examine their communication mechanisms, applications within agriculture and some potential future research directions and opportunities. In addition, a succinct summary is provided in Table VII.
VI-A Introduction to the Internet of Fungus
@@
VI-B Communication Mechanism within Fungal Networks
@@ VI-C Applications in Agriculture
TABLE VII: Applications, Challenges and Research Directions of Internet of Fungus in Smart Agriculture
@@
VI-D Future Directions & Research Opportunities
The IoF is an innovative paradigm that holds immense potential in revolutionizing the agricultural sector. It can improve crop & soil health, boost productivity and provide a sustainable ecosystem. This concept underscores the importance of investigating fungal networks to advance agricultural practices. Some prospective areas of research are mentioned below.
@@ VI-D Future Directions & Research Opportunities
VII Internet of Energy Harvesting Things: Sustainable Power for Agriculture
In the IoT, ensuring sustainable power for a vast network of interconnected devices is a critical challenge. Conventional power sources, such as direct electrical connections or bulky batteries, are not viable for maintaining the multitude of small, ubiquitous smart objects continuously operational. Energy harvesting emerges as a vital solution, enabling these devices to operate autonomously and overcome power constraints. In this section, we will delve into the realm of energy harvesting for smart agriculture, including its techniques, current trends, and future directions & research opportunities. Furthermore, a summary of applications, challenges and future research directions is presented in Table VIII.
TABLE VIII: Applications, Challenges and Research Directions of Internet of Energy in Smart Agriculture
@@ VII Internet of Energy Harvesting Things: Sustainable Power for Agriculture
VII-A Overview of Internet of Energy
VII-B Energy Harvesting Techniques
  1. 1.
@@ VII-B Energy Harvesting Techniques
VII-C Emerging Trends in Energy Harvesting
@@
VII-D Future Directions & Research Opportunities
@@
VIII The Synergy of Internet of Vehicles, Internet of Drones & Internet of Space Things for Efficient Farming
The convergence of IoT, Internet of Vehicles (IoV), Internet of Drones (IoD) and Internet of Space Things (IoST) is anticipated to significantly transform the agricultural sector. This integration will provide farmers and researchers with real-time insights, enhanced resource management, and the facilitation of sustainable operations & practices. Specifically, internet connected tractors, harvesters and other agriculture machinery will enhance the productivity levels. Similarly, IoD and IoST frameworks are set to empower farmers to make data-driven and informed decisions. Fig. 9 depicts the synergistic impact of IoST, IoD, IoV, and IoEn on agriculture. In this section we will discuss the pivotal role of IoV, IoD and IoST for augmenting the productivity of agriculture sector. Additionally, some prospective future directions will also be explored, with a summary provided in Table IX.
VIII-A Internet of Vehicles: Optimizing Agricultural Operations
@@
VIII-B Internet of Drones: A new Horizon in Pest Management and Crop Surveillance
@@
TABLE IX: Applications, Challenges and Research Directions of IoV, IoD & IoST in Smart Agriculture
@@
VIII-C Internet of Space Things: Taking Agricultural Data to New Heights
VIII-D Future Directions & Research Opportunities
@@
IX The Nexus of 6G and Precision Agriculture
The Internet of Things (IoT), a vast network connecting a multitude of objects and devices, has been instrumental in fostering novel services and technological innovations. As the number of IoT devices continues to surge, there is an increasing demand for robust network infrastructure. Alongside the widespread adoption of IoT, several challenges emerge that require resolution. This interconnected ecosystem depends heavily on sophisticated communication technologies, such as 5G and the upcoming 6G network [185], which are crucial for enabling advanced hybrid and multi-cloud solutions. In this section we will discuss the critical role of 6G technology in enhanced connectivity for modern agriculture, key technological enablers of 6G and some prospective future research directions. In addition, a summary of applications, challenges and future research directions is presented in Table X.
IX-A Enhanced Connectivity for Modern Agriculture
@@
IX-B 6G Technological Enablers for Agricultural Advancements
@@
    TABLE X: Applications, Challenges and Research Directions of 6G in Smart Agriculture
    6G
      High energy consumption due to high frequency bands [204].
      Expensive 6G infrastructure.
      Lack of legal & policy frameworks.
      Integrating AI with 6G for NW management in IoAT.
      Zero-Energy sensors for IoAT.
      THz spectroscopy for metal contaminant detection in food.
@@
IX-C Future Directions & Research Opportunities
@@
X Blockchain Technology for Smart & Sustainable Agriculture Sector
Blockchain technology is a digital system that enables electronic documentation, validation and verification thereby eliminating the need for intermediaries. It ensures data and information visibility for all participants, with all records being immutable that are impervious to tampering or erasure. The latest trends and improvements in blockchain technology offer some exciting features that are well suitable for agriculture. This section provides an overview of the blockchain technology, explores its applications & challenges within the agricultural sector, and highlights potential future research directions and opportunities, with a summary presented in Table XI.
X-A Overview of Blockchain Technology
@@
Blockchain’s contributions to high security, enhanced trust and improved efficiency enable its application beyond mere financial transactions. It extends into areas such as supply chain management, voting systems, and data protection, thus reshaping our digital interactions with a new level of transparency and immutability [250]. Recent studies further highlight the role of distributed architectures in blockchain systems for mitigating single-point failures and improving scalability, efficiency, and security—key features for agricultural applications such as fog radio access networks [251] and other scalable frameworks [252].
X-B Blockchain Technology in Agriculture
@@
    Centralized storage, typical of IoT-based systems, makes data susceptible to loss and manipulation. A distributed database can mitigate this issue by facilitating a decentralized storage. Blockchain, in this context, provides a secure, scalable and robust framework for IoT and ICTs in the realm of smart farming. Several smart farming models based on blockchain & IoT have been successfully implemented. In [264] authors introduced an Ethereum blockchain network designed for automating water control in irrigation systems. A generalized blockchain based security architecture for smart sensing in IoT driven agriculture is presented in [265]. [266] explores smart contracts for enabling scalable data sharing in the context of IoT for smart agriculture. Comprehensive literature reviews on blockchain & IoT are provided in [267], [268], [269], [270], [271], underscoring the importance of these technologies in advancing precision agriculture.
@@
    TABLE XI: Applications, Challenges and Research Directions of Blockchain in Smart Agriculture
    Blockchain Technology
    Transparency & Security   Supply chain management.
      Secure data sharing.
      Agriculture finance.
      Interoperability across blockchain networks.
      High energy consumption by blockchains.
      Lack of regulatory policies.
      Addressing scalability issues in large scale deployments.
      Integrating energy harvesting to reduce costs.
      Tokenizing agricultural assets for new opportunities.
      Develop quantum blockchain hybrid models.
      Agricultural insurance to facilitate compensations for losses.
@@
X-C Future Directions & Research Opportunities
@@
XI Machine learning in Precision Agriculture
The incorporation of IoT into agriculture has significantly increased this sector’s efficiency and productivity through the goal of full process automation. There is a growing emphasis among researchers on the use of intelligent computational methods to further augment agricultural practices. Numerous ML algorithms have been developed and applied for purposes like monitoring, controlling, predicting, and smart decision making. ML applications span across various sectors including agriculture, education, healthcare, military, finance, transportation, marketing, manufacturing, and cybersecurity. In this section, we will explore the basics of ML, investigate its utilization in precision agriculture, and identify prospective directions for future research. In addition, a summary of applications, challenges, and future research directions is also presented in Table XII.
@@ XI Machine learning in Precision Agriculture
XI-A Overview of Machine Learning
@@
XI-B Machine Learning Applications in Precision Agriculture
@@
    TABLE XII: Applications, Challenges and Research Directions of Machine Learning in Smart Agriculture
@@
XI-C Future Directions & Research Opportunities
@@
XII Discussion
This section examines key factors for successful IoE adoption in agriculture, offering insights to support the sustainable integration of IoE technologies across diverse agricultural contexts.
XII-A Importance of User-Centric Design for IoE Adoption
The effective integration IoE in agriculture fundamentally relies on a user-centered design approach, particularly to facilitate non-technical users, such as farmers. A CISCO survey indicates that 60% of IoT initiatives encounter adoption challenges when usability and training requirements are overlooked. Addressing these challenges requires attention to the following key aspects as outlined in Table XIII:
  1. 1.
    Understanding End-Users Needs: Direct engagement with end-users / farmers is essential to comprehend their daily challenges, workflows, and specific requirements [331]. This understanding is crucial for developing solutions that are both relevant and beneficial to their needs.
  2. 2.
    Ease of Use and Accessibility: The broad scope of IoE in agriculture, spanning biological, molecular, and ML-driven processes, often produces complex data that may be difficult for farmers to interpret. Simplified dashboards, mobile alerts, and visual indicators (e.g., color codes) will help farmers to make informed decisions. Integrating advanced IoE data (e.g., from molecular sensors and designer phages) into familiar IoT platforms will enhance accessibility and reduces complexity.
    Figure 15: Projected market size for smart agriculture by 2027[332]
    (a) Projected market size for AI in smart agriculture
    (b) Percent usage of AI & IoT in smart agriculture
    Figure 16: AI Market Projections and Usage in Smart Agriculture
  3. 3.
    Affordability: Making IoE solutions more affordable for small and medium-scale farmers is essential for promoting widespread adoption. Governments can support affordability through subsidies, and public-private partnerships can provide accessible financing and deployment models.
  4. 4.
    Localization and Training: Robust IoE implementation in agriculture requires solutions tailored to specific local contexts which addresses language, cultural nuances, and region-specific requirements.
XII-B Quantitative Impact of IoE Technologies on Agricultural Productivity
The integration of specialized IoE domains in agriculture offers substantial benefits across critical areas, such as yield enhancement, water efficiency, and cost reduction. Leveraging advanced IoT, AI, and nano-enabled devices / entities allows agricultural practices to become more data-driven and precise, significantly boosting productivity and promoting sustainability. Table XIV presents an overview of these key benefits and their impact on modern agriculture.
Projections indicate strong growth in the global smart agriculture market, with estimates from the European Agricultural Sector suggesting it will reach USD 34.1 billion by 2026 [332]. Fig.15 further illustrates the projected market size of various smart agriculture applications by 2027, underscoring the increasing demand and adoption of IoE technologies in agriculture.
Fig. 16 illustrate the significant growth and varied applications of AI and IoT within precision agriculture [333]. Fig. 16(a) presents the projected expansion of the global AI in agriculture market, anticipated to reach USD 10.2 billion by 2032. This growth is largely driven by advances in ML, predictive analytics, and computer vision—technologies essential for crop yield forecasting, disease detection, and resource optimization. In Fig. 16(b), the distribution of IoT and AI applications reveals that 35% of usage focuses on farm management activities (e.g., irrigation and crop monitoring), while 20% is dedicated to weather forecasting for informed scheduling, and 15% supports agricultural information systems for training and knowledge dissemination.
These findings highlight the increasing role of IoT and AI in promoting sustainable, efficient, and resilient agricultural practices, underscoring their potential to meet global food security and environmental goals.
TABLE XIII: Key Aspects of User-Centric Design for IoE Adoption in Agriculture
Aspect Objective Solution Impact
Understanding End-User Needs Align IoE solutions with farmers’ workflows Direct farmer engagement Increases relevance and usability
Ease of Use and Accessibility Simplify complex data for non-technical users User-friendly dashboards, alerts, visual indicators Enhances decision-making for farmers
Affordability Make IoE accessible to small/medium-scale farmers Government subsidies, public-private partnerships Expands technology adoption
Localization and Training Adapt IoE to local language and cultural contexts Regional language support, practical training programs Boosts adoption and reduces training time by 30%
Data Security and Privacy Protect sensitive agricultural data Blockchain frameworks, decentralized architectures Strengthens data integrity
TABLE XIV: Quantitative Impact of IoE Technologies in Precision Agriculture
Technology Application Cost Reduction Yield Improvement Water Savings
IoT Crop monitoring, automated irrigation, soil management 15-20% reduction in input and 60% labor costs [334] 20-30% [335] upto 30% [21]
5G/ 6G Real-time remote monitoring 20-40% reduction in operational costs [336] 30-40% (Est. similar tech) 40% [337]
IoD Aerial crop health monitoring, targeted spraying 30% general cost reduction [338] 20% [339] Typically upto 30%, while upto 90% using ULV spraying [340]
IoV Autonomous farming tasks, supplychain, logistics 20% fuel saving [341], 30% op cost [342] 10-15% [343], [342] -
IoST Satellite-based macro monitoring, remote sensing 5% reduction in operational costs [344] 10-12% [345] 100% [346] using private satellite data 5-10% [344]
ML Predictive analytics and forecasting 22% reduction in operational costs [347] 30% [18] 15-30% [18]
Blockchain Supply chain traceability, transactions 30% reduction in transaction costs [348] Indirect, 10-12% reduction in post-harvest losses (Est. similar tech) up to 30% integrating Blockchain & IoT [349]
IoEn Smart energy management, Renewable energy integration 30% reduction in cost [21] 10-15% [350] 20-30% using agrivoltaics [351]
Nanoparticles Targeted delivery upto 50% [352] 25-30% [352] upto 60% through water retention [353]
Designer Phages Disease control 40% [354] Upto 10% [355] Indirect, 10-12% reduction in diseases (Est. similar tech)
IoF Soil health and nutrient sharing 20-40% reduction in fertilizer costs [356] 10-40% [356] [357] 12-17% [358]
XII-C Long-Term Maintenance and Scalability Challenges
While IoE technologies offer substantial benefits, they also introduce challenges related to long-term maintenance and scalability. Key challenges include:
  1. 1.
    Device Maintenance and Durability:IoT devices utilized in agriculture are exposed to harsh environmental conditions, including extreme temperatures, humidity, and dust, which can accelerate wear and tear. Mitigation strategies include:
    • Developing devices with durable materials and protective casings.
    • Implementing proactive maintenance techniques, such as predictive maintenance, to identify and address potential issues before failure. By understanding different types of predictive maintenance solutions, farmers can make more informed decisions, resulting in substantial cost savings.
  2. 2.
    Energy and Connectivity Constraints: Many IoE devices operate in remote agricultural areas where reliable power sources and internet connectivity are limited. For instance, global internet access stands at 81% in urban areas and 50% in rural areas, with rural regions of least developed countries at just 26% [359]. This lack of connectivity poses challenges for real-time data transmission. Suggested solutions include:
    • Utilizing low-power sensors and communication modules to extend battery life.
    • Integrating energy-harvesting technologies to provide continuous, low-cost power.
    • Deploying LPWAN technologies, such as LoRaWAN, which offer extended range and low power consumption, ideal for remote areas.
  3. 3.
    Scalability Across Diverse Farm Types: Adoption of agricultural technology varies significantly by region and farm type. For example, 61% of farmers in Europe and North America plan to adopt at least one technology within the next two years, compared to only 9% in Asia [360]. Larger farms are more likely to adopt these technologies due to greater resources, highlighting the need for scalable solutions suitable for smaller farms. Suggested strategies include:
    • Designing modular IoE systems that can be customized for different farm sizes.
    • Utilizing cloud platforms to handle large volumes of data, ensuring scalability and accessibility.
    • Adhering to industry standards for interoperability between devices and systems, which facilitates easier scaling.
XII-D Data Security and Privacy Challenges
Security is essential in IoE-enabled agriculture due to the sensitive nature of agricultural data. In smart farming systems, vulnerabilities can expose critical data related to crop health, soil conditions, and financial details, impacting both farm management and market dynamics. Studies show that 84% of IoT adopters have experienced security breaches, underscoring the importance of robust security measures [330]. Key risks and mitigations include:
  1. 1.
    Unauthorized Access and Data Breaches: IoT devices often collect sensitive data. Implementing multi-factor authentication, role-based access control, and encryption for data in transit and at rest can prevent unauthorized access.
  2. 2.
    Data Tampering and Integrity Attacks: Attackers can alter IoT sensor data, leading to incorrect farming decisions. Cryptographic hash functions and blockchain technology can ensure data integrity and create immutable records, enhancing traceability
  3. 3.
    Inference Attacks: Even anonymized data can be vulnerable to inference attacks when combined with other data sources. Data aggregation, anonymization, and access controls are essential to safeguard privacy.
  4. 4.
    Denial of Service (DoS) Attacks: DoS attacks on IoT devices can disrupt critical operations like automated irrigation. Network security measures, such as firewalls and intrusion detection systems, along with system redundancy, can mitigate this risk.
  5. 5.
    Supply Chain Vulnerabilities: Compromised components in the supply chain can introduce vulnerabilities. Vendor assessments and secure boot mechanisms can verify the integrity of device firmware and prevent tampering.
A comprehensive strategy that combines technical safeguards, policy development, and continuous monitoring is needed to address these risks. Blockchain-based security frameworks and decentralized architectures are recommended to further enhance data integrity and resilience in IoE-enabled agriculture.
XIII Conclusion
This paper presents an in-depth analysis of agriculture’s role in addressing global challenges such as food security and environmental sustainability. Through a structured research methodology, we examine and emphasize the transformative impact of IoT technologies - including remote sensing, wireless sensor networks, and big data analytics - on smart agriculture, we extend these concepts into the IoE paradigm for further improving the agricultural sector.